OpenStudy (anonymous):

how do I rearrange this equation for q: 16 = 25/q * 80 + (6/10)

OpenStudy (anonymous):

@shannon342 @shkrina

OpenStudy (anonymous):

16-6/10 = 25/q*80 154/10=25/q*80 q=25*10/154*80